OpenStudy (anonymous):

would the limit for lim x-->0 e^x sin x be equal to 0??

OpenStudy (anonymous):

or 1.....?

OpenStudy (psymon):

Just 0. Because yeah, e^0 = 1, but sin(0) = 0, so 0 *1 = 0

OpenStudy (anonymous):

thought so :) so you can basically just plug in the number to the equation ?

OpenStudy (anonymous):

lol sorry for dumb questions, i'm not very bright :(

OpenStudy (psymon):

Yep. As long as you dont get an undefined answer. sin(0)e^0 isnt undefined, so I could care less if it still spits out 0 xD

OpenStudy (anonymous):

what do you do when it asks about ints?? for ex lim x--> 7/2- int (2x-1) ? :o

OpenStudy (psymon):

int?

OpenStudy (anonymous):

that's the hw problem >__< D:

OpenStudy (anonymous):

what does that even mean? lol

OpenStudy (psymon):

Not sure what int would be, but wouldnt it just be 6? xD

OpenStudy (anonymous):

that would be my guess as well, but i graphed it and i get like 4.5 when i trace it

OpenStudy (anonymous):

?D:

OpenStudy (psymon):

In order for you to graph it wouldnt you have to know what int is? xD

OpenStudy (anonymous):

ok, we'll just go with 6(: jaja

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

QUESTIONNNN

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

when your looking at limits approaching infinity

OpenStudy (anonymous):

do you look at what is going to left/ right of the graph or whats going up and down??

OpenStudy (psymon):

Well, limits are literally "what is the y-value at x = ?" So youre checking left and right to observe up and down. So for infinite limits, you check values appraoching from the left and appraoching from theright and see if they both shoot off to infinity or both shoot off to negative infinity. If they go in opposite directions then the limit doesnt exist.

OpenStudy (anonymous):

so correct me if i'm wrong but x-->infinity (x+sin(x))/(x+cos(x)) is DNE?

OpenStudy (psymon):

In this case, no actually.

OpenStudy (anonymous):

??:( whyyyyyyyy? lol

OpenStudy (psymon):

Im checking xD I just punched into my calculator real fast without checking it, but Im working on confirming.

OpenStudy (psymon):

Okay, yeah, we have to do it this way. Have you see the method of dividing everything by the highest power in the denominator?

OpenStudy (anonymous):

.__. no lol

OpenStudy (psymon):

Ah x_x Okay, so a normal trig when a function has x -> infinity or negative infinity is to find the highest power in the denominator and divide all terms in the function by that power. So in the denominator, all we have is a 1st power of x. So we want to divide all terms by x to start: \[\frac{ \frac{ x }{ x }+\frac{ sinx }{ x } }{ \frac{ x }{ x }+\frac{ cosx }{ x} }\]which simplifies to: \[\frac{ 1+\frac{ sinx }{ x } }{ 1+\frac{ cosx }{ x } } \]Now we have to take x to infinity. Now, two things to mention. In general, if we have a function like: \[\lim_{x \rightarrow \infty}\frac{ x^{a} }{ x^{b} }\] and a > b, then the limit is infinity. If b > a, then the limit is 0. So thats something to be aware of. However, when you have trig functions you have to be a bit more careful. You just just say it goes to 0 because theres an x in the denominator. Now, some textbooks consider these identities, some make you show them, but: \[\lim_{x \rightarrow \infty}\frac{ sinx }{ x }= 0\] \[\lim_{x \rightarrow 0}\frac{ cosx }{ x }=0\]Now, if thse werent identities and you had to prove them, youd have to prove them with squeeze theorem. Basically, the squeeze theorem requires you to show, using inequalities, that the left inequality and the right inequality are the same as limit goes to infinity. So this is how itd bedone: \[-1 \le sinx \le 1\]The range of the sine function is -1 to 1, so this inequality is valid. But we have sinx/x, so lets divide everything by x \[\frac{ -1 }{ x } \le \frac{ sinx }{ x } \le \frac{ 1 }{ x } \]Now as I said, the goal is to have the left inequality and the right inequality to have the same limit as x goes to infinity. Well, as was just mentioned above, if the power of x is higher in the denominator, which it is in -1/x and 1/x, then limit is 0. So basically, since the left inequality is 0 and the right inequality is 0, sinx/x also has to be 0. This exact same thing can be done with cosx/x. Either way, we end up with: \[\frac{ 1+0 }{ 1+0 }= 1\] I know thats a lot of info, so feel free to ask anything lol.

OpenStudy (anonymous):

YOU. ARE. AWESOME. omg. thank you so much!!!!

OpenStudy (anonymous):

i get it :)

OpenStudy (psymon):

Okay, cool xD Just yeah, if you ever come up with a trig function with these limits, it may require yout o try and mess around with the squeeze theorem, especially with ones that go to infinity.

OpenStudy (anonymous):

got it. :)

OpenStudy (anonymous):

seriosuly thanks so much for that

OpenStudy (anonymous):

i would give you a thousand medal just for that long long explanation lol

OpenStudy (psymon):

Yah, np, lol. You wont have to be doing much more with limits once you finish that chapter, but squeeze theorem and knowing about dividing everything by the highest power in the DENOMINATOR or checking limits left and right. All the techniques you can use xD

OpenStudy (anonymous):

when being asked question like at what point is the discontinuity, you're looking for y value right?(:

OpenStudy (anonymous):

lol you're a life saver:D

OpenStudy (psymon):

Well, depends. It may be wanting the whole coordinate. Its either going to be the whole coordinate or just the x-coordinate, theyre not going to be asking the y-value if it just asks wheres the discontinuity.

OpenStudy (psymon):

Most like x-coordinate really, since an asymptote we would just name as x = ?

OpenStudy (anonymous):

what does it mean when it says find a power function end behavior model? :o

OpenStudy (psymon):

That sounds weird, lol. Power function? O.o

OpenStudy (anonymous):

yeahh .__. lol

OpenStudy (anonymous):

i have no clue lol

OpenStudy (psymon):

Yeah, no idea. End behavior is fine on its own, direction the graphs starts from and the direction the graph ends towards. Power function, dunno.

OpenStudy (anonymous):

so when it says find a right end behavior model and a left end behavior model of a function you just describe, draw it.........? .__.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

f(x)=ln abso value(x) +sin x

OpenStudy (anonymous):

JK don't have to study that one :)

OpenStudy (anonymous):

ok, i should REALLY stop bugging you now lol!

OpenStudy (anonymous):

you've been tooooons of help thanks yo! (:

OpenStudy (anonymous):

seriously. you are awesome!

OpenStudy (psymon):

Lol, its fine xD Sorry I didnt respond immediately xD

OpenStudy (anonymous):

no seriously thank you so much :)

OpenStudy (anonymous):

if it's fine i have just one last question for you and then we can part ways :) lol

OpenStudy (psymon):

Go for it.

OpenStudy (anonymous):

what would the graph look like if lim f(x) did not exist and lim f(x)=f(2)=3 ? as x->2

OpenStudy (psymon):

Sounds like its basically saying: \[\lim_{x \rightarrow 2}f(x) = 3\] But its just written in an odd way O.o

OpenStudy (anonymous):

hm, so it'd be 3?

OpenStudy (anonymous):

i mean sorry lol

OpenStudy (anonymous):

that was completely wrong

OpenStudy (anonymous):

wouldn't it be a discont graph that has jump disc?

OpenStudy (psymon):

I guess im not sure how else to interpret it. Is it written exactly like that?

OpenStudy (anonymous):

yeaahh

OpenStudy (psymon):

Yah, itd have to be a jump discontinuity it looks like. Only way for f(2) to be defined but not have a limit.

OpenStudy (anonymous):

ok, i'll just ask my teach to see if my sketch is correct :)

OpenStudy (psymon):

I think the qustion is too vague to have a perfect sketch, just as long as you have some sort of graph that has a coordinate (2,3) that exists but then jumps to some other y-value ax x = 2 xD

OpenStudy (anonymous):

alrighty, thank you @Psymon for the very useful help!

OpenStudy (psymon):

np ^_^

OpenStudy (anonymous):

yeah :) i got that lol

OpenStudy (anonymous):

THANK YOU.

OpenStudy (anonymous):

bye! :)

OpenStudy (psymon):

bye :3